### v1

@@ 3 System architecture
  • Multi-view text generator. Based on the features of three views extracted from the previous module, a novel multi-view prompt engineering method is proposed, which aims to guide the LLM to generate the texts of descriptions and summaries across different views accurately, efficiently, and cost-effectively. To achieve the above goals, function description prompt template and view summary prompt template are designed to provide a unified paradigm for LLM to generate view texts, thus facilitating LLM in generating view texts for a given APK in a more standardized way. (Refer to Section 4.2 for further details.)
  • Detection classifier. Given the texts (descriptions and summaries) from different views, this module transforms all the texts into the machine-readable representation vectors, which are then concatenated into a single representation vector for describing the behavioral semantic information of the APK. Then, a DNN-based classifier is developed to learn the potential importance of representations and give its prediction (i.e., a given unknown APK will be predicted to be malicious or not). (Refer to Section 4.3 for further details.)
  • Diagnostic report generator. To gain a more intuitive understanding of the potential malicious behaviors of a given APK, rather than merely obtaining a binary result (i.e., malicious or benign), this module inputs the descriptions and summaries from different views, along with the detection results, into the specifically designed diagnostic report prompt template. The LLM is then employed to generate a diagnostic report for the given APK, which offers preliminary insight into potential behaviors and provides a foundation for further exploration and validation. (Refer to Section 4.4 for further details.)
@@ 4.1 Feature selection and extraction
To describe the behavioral semantic information of Android applications in a more comprehensive way, inspired by Drebin (Arp et al. (2014)), a classical work based on feature engineering in Android malware detection, we select four main feature types, namely, permission, API, URL, and uses-feature. Referring to Drebin, we subdivide permission into requested permission and used permission, as well as subdivide API into restricted API and suspicious API to further characterize the relevant behaviors. These features are then organized in the form of views to further model the behavioral semantics between them. Note that since a large number of applications do not hardcode URL and declare uses-feature, their number and frequency are much smaller than that of permission and API. Therefore, dividing them into separate views leads to unnecessary resource consumption and also renders a highly unbalanced information content expressed between the views. On the other hand, combining URL and uses-feature into one view facilitates formatting uniformity across all views. Based on the above considerations, the features are organized into three views, i.e., Permission View, API view, and URL & use-feature View. The detailed types of features and views, as well as their descriptions, are shown in Table 1.
@@
To extract these features, we employ Androguard (Desnos and Gueguen (2018)) for decompiling the APK file. Then, a static analysis based extractor is developed to automatically identify and extract relevant features from the AndroidManifest.xml file and class.dex file. Notably, we utilize Pscout (Au et al. (2012)) to obtain used permission and restricted API, which compose the mapping relationship.
@@ 4.2 Multi-view text generation
To describe the content within the Permission View, API View, and URL & uses-feature View, we employ LLM to serve as a domain expert in generating descriptions and summaries, which not only enables sufficient reasoning and summarization of potential behaviors from these views, but also further explores the explicit semantic of each view. In order to fully utilize and unleash the power of LLM in the Android domain, as well as to ensure it outputs what we need in a uniform and standardized format, meticulous prompt engineering is critical. It is worth pointing out that the LLM employed in our work is gpt-4-1106-preview (OpenAI (2024)), which stands out as one of the most well-known and capable model in its field.
In summary, we design a multi-view prompt engineering method as shown in Fig. 2 to generate descriptions and summaries, which fulfill the specific requirements of each view. Specifically, the text generation task for each view is divided into two distinct phases: function description generation and view summary generation. For these phases, we designed function description prompt template and view summary prompt template, enabling the LLM to produce detailed descriptions and summaries for each view.
@@ 4.2.1 Function description generation
For a given APK, the feature extractor extracts the feature list based on different feature subtypes. By sequentially injecting each specific feature into function description prompt template and subsequently inputting it into the LLM to generate the corresponding function description, we obtain a set of feature function key-value pairs for each feature subtype (i.e., function description list). This list delineates all specific features and their respective functions within each feature subtype, serving as a preliminary overview of the potential behavior. The function description list on the one hand, is used as one of the important sources for classification detection, and on the other hand, it will be used as an input for obtaining view summary, which is a further summary of the behavior of the whole view.
It is imperative to utilize the LLM for generating function descriptions, rather than extracting descriptions directly from official documentation (Android Developers (2024)). This necessity arises primarily for two reasons: 1) Official documents contain various dimensions for system-level permissions and APIs, complicating the automatic and direct extraction of pertinent information for function descriptions. 2) For user-defined permissions, APIs, numerous URLs and other data, the LLM possesses robust summarization capabilities that surpass any alternative automated search and summarization methods. For instance, the permission "com.google.android.c2dm.permission.RECEIVE" cannot be directly located and summarized from Android’s official documentation, however, through carefully crafted prompt engineering with the LLM, we can obtain its associated function description, namely, "allows receiving push notifications from Google Cloud Messaging (GCM)".
@@ 2 Example {Feature type}: {Example feature}
System setup + Example ×3 + Input
@@
  • Example. Three carefully crafted examples of specific generative patterns and styles are subsequently provided for LLM to assimilate.
@@ 4.2.2 View summary generation
In order to explore the deeper potential behavioral information hidden under the features and their functions, we employ LLM to further generate summary for each view by using the function descriptions list generated by the aforementioned modules. Specifically, in this sub-module, we try to construct a step-by-step template to inform LLM in detail about the steps and requirements for generating view summary as well as the explanations of the input information, which ensures that LLM can analyze and deduce according to the criteria, thus producing consistently formatted summaries.
@@ 4 Nouns Interpretation <Nouns Interpretation>:
Injecting the pertinent function descriptions lists into the prompt structured from the view summary prompt template enables the LLM to be guided to generate an appropriate view summary. This facilitates further reasoning and summarization based on the features and their functions to obtain more comprehensive behavioral insights. In summary, we provide the LLM with a detailed chain-of-thought reasoning to empower it to generalize and summarize the behavior of the views with the same thought process, which not only enhances the training of classification models but also aids in producing more insightful diagnostic reports.
@@ 4.2.4 Multi-view prompt implementation
Decomposing the text generation task into two sub-phases (i.e., function description generation and view summary generation) allows us to sequentially connect the entire process in a chain-like manner, yielding higher-quality descriptions and summaries. Besides, considering that our task contains multiple views, it is essential that the function description prompt template and view summary prompt template are able to adapt to multi-view, thus outputting relevant descriptions and summaries aligned with the diverse contents within these views. Consequently, our multi-view prompt establishes a unified approach to integrating the contents from various views into the templates based on formatting guidelines, facilitating the generation of descriptions and summaries pertinent to each view. Ultimately, through multi-view prompt engineering, for each view, this module leverages the LLM to generate two function descriptions and one view summary. Based on the generated text, detection tasks can be further carried to generate diagnostic reports.
@@ 25       Generate the label $f_{n}$ using trained MLP;
To train the classifier, the initial task of this module includes converting the texts from function descriptions and view summaries generated by the LLM into a machine-readable vector format. Considering the need for consistency and recognizing that text embedding is not the focused innovation in this work, we employed OpenAI’s embedding model text-embedding-ada-002 (OpenAI (2024)) to transform the generated descriptions and summaries from different views into 1536-dimensional representation vectors. These vectors are then concatenated into a single vector encapsulating the comprehensive behavioral semantic information of the APK from all three views.
Utilizing our dataset and the representation vectors of APKs, a Deep Neural Network (DNN)-based classification model is trained, as detailed in Section 5.4 about the selection of the classification model. Whenever the detection task of an unknown APK is performed, after obtaining the representation vector of that APK through the aforementioned steps, it can be directly fed into the trained classifier to determine whether the application is malicious or not.
@@ 4.4 Diagnostic report generation
In practical detection environments, merely obtaining a binary detection outcome (malicious or benign) is often insufficient. It is crucial to provide a diagnostic report for unknown APKs, which can identify potentially malicious behaviors and guide further investigation or detection. Existing feature engineering based machine learning or deep learning methods still face significant challenges in generating readable and instructive diagnostic reports.
@@ 5 Experiment and evaluation
  • RQ3: How does AppPoet select a classification model to ensure effective detection?
  • RQ4: Are the diagnostic reports generated by AppPoet instructive and valid?
@@ 5.1 Experiment setup
Dataset. To objectively assess the multifaceted performance of AppPoet, our dataset comprises 11,189 benign Apps and 12,128 malicious Apps sourced from AndroZoo (Allix et al. (2016)), a dataset collects Apps primarily from official App stores like Google Play and uses VirusTotal (VirusTotal (2024)) to determine the nature of each App.
@@ 5.2 RQ1: Performance of AppPoet
In order to explore the detection capability of AppPoet in real-world applications, Drebin (Arp et al. (2014)) has been selected as the baseline method for comparison in this section, which is based on the real-world application dataset outlined in Section 5.1. Drebin is a representative work based on feature engineering to Android malware detection task, which extensively captures a wide range of features and utilizes one-hot encoding to obtain representations. A classifier is then trained for the detection task. To ensure consistency, inputs for Drebin are selected to match the feature subtypes of AppPoet shown in Table 1. In addition, to enhance Drebin’s performance, we train an MLP version of the classification model as Drebin’s variant. For our method, AppPoet parses the feature subtypes according to the method of multi-view prompt proposed in Section 4.2 and uses the method of Section 4.3 to train the classification model. We randomly divide 80% from the real-world application dataset for training and the rest for testing.
Figure 3: Comparison of malware detection performance for different methods.
Fig. 3 presents the detection results of different methods. From the results, it is evident that AppPoet outperforms both Drebin-SVM and Drebin-MLP across all indicators. This experiment demonstrates that representation vectors based on feature engineering have already possessed strong representational capabilities. Drebin-SVM has already achieved excellent detection performance, which is further enhanced by utilizing the more advanced MLP classification model. However, Drebin’s representation vectors are limited to representing information about the features themselves and lack explicit mining of function and behavioral semantics among the features. In our method, the semantic richness of the representation vectors is significantly enhanced by leveraging the powerful reasoning and summarization capabilities of LLM to generate function descriptions and view summaries of known facts. Moreover, from the results, it can be concluded that our method yields better performance in real detection tasks, evidencing a significant improvement in the semantic richness of representation vectors.
@@ 5.3 RQ2: Performance of prompt engineering
Since AppPoet’s detection relies on the descriptions and summaries generated by LLM, the quality of these texts significantly influences the representational capabilities of the vectors, and consequently the detection outcomes. Therefore, discussing the effectiveness of the multi-view prompt engineering method proposed in this paper is imperative. In this section, a series of ablation experiments are conducted to assess the impact of view, description and summary on the classification outcomes. Then, the ablation experiments about prompt design are conducted to assess the validity of our method’s workflow. Finally, the utility of the memory component in AppPoet is evaluated, which is developed to enhance efficiency and minimize costs.
@@
Ablation experiments about workflow design. Our multi-view prompt engineering approach is to generate corresponding function descriptions and view summaries for different views in a multi-view, multi-phase manner. With this batch-phase design, we can enhance LLM’s attention to each detail and generate textual information that is as objective and adequate as possible, which not only ensures the model’s detection performance, but also beneficial for improving the richness of diagnostic reports. In order to validate the effectiveness of our approach, we designed the prompt template with multi-view, no phase, and the prompt template with multi-phase, no view (see the Appendix A for details). Specifically, the multi-view, no phase prompt engineering designs a prompt template for each of the Permission, API and URL & uses-feature Views. After injecting a list of features extracted from the static extractor into it, the LLM generates the function description and view summary of the view’s features directly based on the template’s chain guidance. Multi-phase, no view prompt engineering, in contrast, divides the generation task into two phases, description generation and summary generation, designs a prompt template for each of these phases, and follows the AppPoet approach of in-context learning and thought of chain guidance approach. But unlike AppPoet, in this prompt engineering method, we converge all views together and output both the function description and view summary for all three views at once. To save experimental overhead while validating the performance of the method, we randomly sampled 1000 malicious samples and 1000 benign samples from the previous dataset for small-scale experiments. Our experiments evaluate the success number of different methods for outputting text and the success number of outputting text that conforms to the required format and can be parsed and embedded by automation. Finally, the samples are divided into 80% training set and 20% test set and the accuracy of the detection is evaluated by the same MLP model. The experimental results are shown in Table 8.
@@
multi phase, no view -   998 | 990   998 | 990 91.96 91.88
no phase, multi view Permission   989 | 991   955 | 960 90.63 90.22
API   997 | 986   907 | 871
URL & uses-feature 1000 | 999 1000 | 999
@@
multi phase, multi $\begin{array}[]{c}\text{multi phase, multi view}\\ (\text{ours})\end{array}$ Permission   1000 | 1000   1000 | 1000 94.50 94.21
API   1000 | 1000   1000 | 1000
URL & uses-feature   1000 | 1000   1000 | 1000
@@
As can be seen from the table, the multi-view, multi-phase approach adopted by AppPoet outputs all the text successfully, and all the text is formatted, parsed and embedding, which obtains a detection accuracy of 94.50% of the ACC and 94.21% of the F1 value. In contrast, neither the multi-view-only, nor the multi-phase-only prompt engineering approach is capable of outputting the full amount of textual information. This is mainly because our experiments are based on utilizing OpenAI’s API to communicate with a remote server network. In the case of batch calls, if the prompt of a single input is too long or too much output is requested, there is a certain probability that the interaction fails due to the network problem of not being able to deliver the message properly. More seriously, the no-phase approach, due to the excessive requirements on the content of a single output from the LLM, can easily lead to the inability of the LLM to output the content in strict accordance with the preset format and requirements, as well as the inability of extracting the effective information to be transformed into embedding vectors for the discriminative process. The results of this method have a large amount of output text that cannot be utilized and transformed. In addition, the detection performance ultimately obtained by these two methods lags significantly behind that used by AppPoet. In summary, the prompt engineering approach we designed can obtain better detection performance while ensuring correct output and transformation.
@@
Memory experiments. In order to evaluate the performance of our designed memory component, we perform a set of comparison experiments. Based on the 2000 samples randomly selected in the previous dataset, we conduct two sets of experiments with memory and without memory respectively while the multi-view text generator is working, and count the average prompt token consumption and response token consumption of each application in the experiments, as well as the the average time taken to generate text. The difference between these two sets of experiments is only whether the memory component is used to store functions. The results are shown in Fig 4.

### v3

@@ 3 System architecture
  • Multi-view text generator. Building on the features extracted from the three views in the previous module, we propose a novel multi-view prompt engineering method. This approach aims to guide the LLM in generating descriptions and summaries across different views. To achieve this, we design function description prompt template and view summary prompt template, providing a unified framework that allows the LLM to generate standardized texts for a given APK. (Refer to Section 4.2 for further details.)
  • Detection classifier. Given the texts (descriptions and summaries) from different views, this module transforms all the texts into the machine-readable representation vectors, which are then concatenated into a single representation vector for describing the behavioral semantic information of the APK. Then, a DNN-based classifier is developed to learn the potential importance of the representations and give its prediction (i.e., a given unknown APK will be predicted to be malicious or not). (Refer to Section 4.3 for further details.)
  • Diagnostic report generator. To provide a more intuitive understanding of a given APK’s potential malicious behavior, this module goes beyond a simple binary result (malicious or benign). It combines the descriptions and summaries from different views, along with the detection results, into a specially designed diagnostic report prompt template. The LLM is then employed to generate a diagnostic report for the given APK, which offers preliminary insight into potential behaviors and provides a foundation for further exploration and validation. (Refer to Section 4.4 for further details.)
@@ 4.1 Feature selection and extraction
To describe the behavioral semantic information of Android applications in a more comprehensive way, inspired by Drebin (Arp et al., 2014), a classical String-based work in Android malware detection, we select four main feature types, namely, permission, API, URL, and uses-feature. Referring to Drebin, we subdivide permission into requested permission and used permission, as well as subdivide API into restricted API and suspicious API to further characterize the relevant behaviors. These features are then organized in the form of views to further model the behavioral semantics between them. Note that since a large number of applications do not hardcode URL and declare uses-feature, their number and frequency are much smaller than that of permission and API. Therefore, dividing them into separate views leads to unnecessary resource consumption and also renders a highly unbalanced information content expressed between the views. On the other hand, combining URL and uses-feature into one view facilitates formatting uniformity across all views. Based on the above considerations, the features are organized into three views, i.e., Permission View, API View, and URL & uses-feature View. The detailed types of features and views, as well as their descriptions, are shown in Table 1.
@@
To extract these features, we employ Androguard (Desnos and Gueguen, 2018) for decompiling the APK file. Then, a static analysis based extractor is developed to automatically identify and extract relevant features from the AndroidManifest.xml file and class.dex file. Notably, we utilize PScout (Au et al., 2012) to obtain used permission and restricted API, which compose the mapping relationship.
@@ 4.2 Multi-view text generation
To describe the content within the Permission View, API View, and URL & uses-feature View, we leverage LLM as a domain expert to generate descriptions and summaries. This approach not only facilitates thorough reasoning and summarization of potential behaviors from these views, but also delves deeper into the explicit semantics of each view. It should be noted that the detection performance of our system depends heavily on the text quality of the LLM outputs. How to guide LLM to fully utilize abilities in order to obtain high quality text possible is an important issue we must consider. In order to fully utilize and unleash the power of LLM in the Android domain, as well as to ensure it outputs what we need in a uniform and standardized format, meticulous prompt engineering is critical.
In summary, we design a multi-view prompt engineering method as shown in Fig. 2 to generate descriptions and summaries, which fulfill the specific requirements of each view. Specifically, the text generation task for each view is divided into two distinct phases: function description generation and view summary generation. For these phases, we designed function description prompt template and view summary prompt template, enabling the LLM to produce detailed descriptions and summaries for each view. In this way, we resolve the text generation task into finer-grained phases and views in the form of workflows, ensuring that the LLM generates the appropriate content step by step. By utilizing this systematic approach, our method achieves impressive detection performance, as demonstrated in Section 5.3, which also reflects the high quality of the generated text.
It is worth noting that the capabilities of the LLM itself are also an important factor in the quality of the text output. The LLM employed in our work is gpt-4-1106-preview (OpenAI, 2024), which stands out as one of the most well-known and capable model in its field. In this way, the lower bound on the quality of the text output can be guaranteed.
@@ 4.2.1 Function description generation
For a given APK, the feature extractor extracts the feature string sequence based on different feature subtypes. By sequentially injecting each specific feature into function description prompt template and subsequently inputting it into the LLM to generate the corresponding function description, we obtain a set of feature function key-value pairs for each feature subtype (i.e., function description list). This list delineates all specific features and their respective functions within each feature subtype, serving as an intuitive natural language overview of the potential behavior. The function description list on the one hand, is used as one of the important sources for classification detection, and on the other hand, it will be used as an input for obtaining view summary, which is a further summary of the behavior of the whole view.
It is imperative to utilize the LLM for generating function descriptions, rather than extracting descriptions directly from official documentation (Android Developers, 2024). This necessity arises primarily for two reasons: 1) The official documents provide extensive information on system-level permissions and APIs, making it difficult to automatically extract concise and accurate function descriptions. 2) For user-defined permissions, APIs, numerous URLs and other data, the LLM possesses robust summarization capabilities that surpass any alternative automated search and summarization methods. For instance, the permission "com.google.android.c2dm.permission.RECEIVE" cannot be directly located and summarized from Android’s official documentation. However, through carefully crafted prompt engineering with the LLM, we can obtain its associated function description, namely, "allows receiving push notifications from Google Cloud Messaging (GCM)".
@@ 2 Example {Feature type}: {Example feature}
System setup + Example ×$k$ + Input
@@
  • Example. Subsequently, carefully crafted examples of specific generative patterns and styles are provided for LLM to assimilate. It is worth clarifying that detailed selection and evaluation of the number $k$ of examples is described in Section 5.3.
@@ 4.2.2 View summary generation
In order to explore the deeper potential behavioral information hidden under the features and their functions, we employ LLM to further generate summary for each view by using the function descriptions list generated by the aforementioned modules. In this sub-module, we create a step-by-step template to guide the LLM on how to generate view summaries. This approach ensures that the LLM can analyze and deduce based on the specified criteria, producing consistently formatted summaries.
@@ 4 Nouns Interpretation <Nouns Interpretation>:
Injecting the pertinent function descriptions lists into the prompt structured from the view summary prompt template enables the LLM to be guided to generate an appropriate view summary. This facilitates further reasoning and summarization based on the features and their functions to obtain more comprehensive behavioral insights. In summary, we provide the LLM with detailed chain-of-thought reasoning, enabling it to summarize the behavior of the views systematically. This approach not only enhances the training of classification models but also helps produce more insightful diagnostic reports.
@@ 4.2.4 Multi-view prompt implementation
Decomposing the text generation task into two sub-phases (i.e., function description generation and view summary generation) allows us to sequentially connect the entire process in a chain-like manner, yielding higher-quality descriptions and summaries. Additionally, since our task involves multiple views, it is crucial that the function description prompt template and view summary prompt template can adapt to this multi-view structure, ensuring that the generated descriptions and summaries align with the diverse content of each view. Consequently, our multi-view prompt establishes a unified approach to integrating the contents from various views into the templates based on formatting guidelines, facilitating the generation of descriptions and summaries pertinent to each view. Ultimately, through multi-view prompt engineering, for each view, this module leverages the LLM to generate function descriptions and view summary. Based on the generated text, detection tasks can be further carried to generate diagnostic reports.
@@ 25       Generate the label $f_{n}$ using trained MLP;
To train the classifier, the initial task of this module includes converting the texts from function descriptions and view summaries generated by the LLM into a machine-readable vector format. Considering the need for consistency and recognizing that text embedding is not the focused innovation in this work, we employed OpenAI’s embedding model text-embedding-ada-002 (OpenAI, 2024) to transform the generated descriptions and summaries from different views into 1536-dimensional representation vectors. These vectors are then concatenated into a single vector encapsulating the comprehensive behavioral semantic information of the APK from all three views.
Utilizing our dataset and the representation vectors of APKs, a DNN-based classification model is trained, as detailed in Section 5.1 about the selection of the classification model. Whenever the detection task of an unknown APK is performed, after obtaining the representation vector of that APK through the aforementioned steps, it can be directly fed into the trained classifier to determine whether the application is malicious or not.
@@ 4.4 Diagnostic report generation
In practical detection environments, merely obtaining a binary detection outcome (malicious or benign) is often insufficient. It is crucial to provide a diagnostic report for unknown APKs, which can identify potentially malicious behaviors and guide further investigation or detection. Existing learning-based methods still face significant challenges in generating readable and instructive diagnostic reports.
@@ 5 Experiment and evaluation
  • RQ3: Are the diagnostic reports generated by AppPoet instructive and valid?
@@ 5.1 Experiment setup
Dataset. To objectively assess the multifaceted performance of AppPoet, our dataset comprises 11,189 benign Apps and 12,128 malicious Apps sourced from AndroZoo (Allix et al., 2016), a dataset collects Apps primarily from official App stores like Google Play and uses VirusTotal (VirusTotal, 2024) to determine the nature of each App.
@@
Classification model selection. Although the design of classification model is not the main focus of this paper, selecting an appropriate model is crucial for achieving accurate detection results. To this end, we evaluate several common models, including CNN (Simonyan and Zisserman, 2014), TextCNN (Chen, 2015), RNN (Elman, 1990), LSTM (Hochreiter and Schmidhuber, 1997), and MLP (Schmidhuber, 2015), using the real-world application dataset collected above. This dataset was split into 80% for training and 20% for testing. Based on the results in Table 7, the MLP model demonstrated the best overall performance. Given its strong ability to preserve the semantic richness of the feature representations, we chose MLP as the classification model for all subsequent experiments.
Table 7: Comparison of different classification model.
Method ACC(%) Precision(%) Recall(%) F1(%)
CNN 95.79 96.04 95.71 95.87
TextCNN 96.03 96.05 96.17 96.11
RNN 95.94 96.92 95.08 95.99
LSTM 96.07 96.68 95.59 96.13
MLP 97.15 97.03 97.39 97.21
@@ 5.2 RQ1: Performance of AppPoet
To evaluate the detection capability of AppPoet in real-world applications, we compare it with several learning-based methods: (1) Drebin (Arp et al., 2014) which is String-based method; (2) LBDB (Sun et al., 2021) which is Image-based method; and (3) MaMaDroid (Onwuzurike et al., 2019) and Malscan (Wu et al., 2019) which are both Graph-based methods. The setup of these baseline methods is explained as follows. For Drebin, to ensure consistency, we align its input features with the feature subtypes used by AppPoet, as shown in Table 1. Additionally, to enhance Drebin’s performance, we train an MLP variant of its classification model. For MaMaDroid, it has two versions that abstract API calls into either family calls or package calls. We configure and conduct experiments on both versions accordingly. For LBDB and Malscan, we conduct experiments and configurations according to their open-source code. For AppPoet, we parse the feature subtypes according to the method of multi-view prompt proposed in Section 4.2 and use the method of Section 4.3 to train the classification model. We randomly divide 80% from the real-world application dataset for training and the rest for testing.
Table 8: Comparison of malware detection performance for different methods
Method ACC(%) Precision(%) Recall(%) F1(%)
Drebin-SVM 94.76 94.60 95.33 94.96
Drebin-MLP 96.06 96.47 96.04 96.26
LBDB 91.39 92.55 88.52 91.27
MaMaDroid-fml 94.86 93.80 96.59 95.18
MaMaDroid-pkg 95.35 94.72 96.51 95.61
Malscan 95.65 95.54 96.15 95.84
AppPoet 97.15 97.03 97.39 97.21
Table 8 presents the detection results of different methods. The results show that AppPoet outperforms all other baseline methods across various metrics. From the results, we can draw the following conclusions: (1) Image-based methods, which directly convert file-level features into images, tend to overlook key semantic information, resulting in the lowest detection performance among the methods. (2) Although String-based methods have limited ability to capture deep semantic information compared to Graph-based methods, their extensive feature extraction allows them to achieve detection performance comparable to that of Graph-based methods. (3) Our method leverages the LLM’s strong reasoning and summarization capabilities to further explore the behavioral semantics of the features, leading to superior performance.
@@ 5.3 RQ2: Performance of prompt engineering
Since AppPoet’s detection relies on the descriptions and summaries generated by LLM, the quality of these texts significantly influences the representational capabilities of the vectors, and consequently the detection outcomes. Therefore, it is crucial to evaluate the effectiveness of the multi-view prompt engineering method proposed in this paper. In this section, we first perform ablation experiments to assess the importance of different views and text descriptions in influencing detection performance. Next, we conduct ablation experiments on workflow design to evaluate the effectiveness of our prompt workflow in generating descriptions and summaries. Following this, selection experiments are carried out to determine the optimal number $k$ of examples required for function description prompt template, balancing the trade-off between quality and efficiency. Lastly, we perform memory and efficiency experiments, where we assess the role of the memory component and evaluate the real-world detection efficiency of AppPoet.
@@
Ablation experiments about workflow design. Our multi-view prompt engineering approach is to generate corresponding function descriptions and view summaries for different views in a multi-view, multi-phase manner. With this batch-phase design, we can enhance LLM’s attention to each detail and generate textual information that is as objective and adequate as possible, which not only ensures the model’s detection performance, but also beneficial for improving the richness of diagnostic reports. In order to validate the effectiveness of our approach, we designed the prompt template with multi-view, no phase, and the prompt template with multi-phase, no view (see the Appendix A for details). Specifically, the multi-view, no phase prompt engineering designs a prompt template for each of the Permission, API and URL & uses-feature Views. After injecting a list of features extracted from the static extractor into it, the LLM generates the function description and view summary of the view’s features directly based on the template’s chain guidance. Multi-phase, no view prompt engineering, in contrast, divides the generation task into two phases, description generation and summary generation, designs a prompt template for each of these phases, and follows the AppPoet approach of in-context learning and thought of chain guidance approach. But unlike AppPoet, in this prompt engineering method, we converge all views together and output both the function description and view summary for all three views at once. To reduce experimental overhead while ensuring a thorough validation of the method’s performance, we randomly select 2,000 malicious samples and 2,000 benign samples from the dataset, forming a balanced subset of 4,000 samples fro small-scale experiments. Our experiments evaluate the success number of different methods for outputting text and the success number of outputting text that conforms to the required format and can be parsed and embedded by automation. Finally, the samples are divided into 80% training set and 20% test set and the accuracy of the detection is evaluated by the same MLP model. The experimental results are shown in Table 10.
@@
multi phase, no view -   1979 | 1984   1977 | 1984 93.44 92.25
no phase, multi view Permission   1962 | 1989   1901 | 1938 91.36 91.31
API   1983 | 1971   1887 | 1910
URL & uses-feature   1990 | 1999   1989 | 1999
@@
multi phase, multi $\begin{array}[]{c}\text{multi phase, multi view}\\ (\text{ours})\end{array}$ Permission   2000 | 2000   2000 | 2000 95.50 95.37
API   2000 | 2000   2000 | 2000
URL & uses-feature   2000 | 2000   2000 | 2000
@@
As can be seen from the table, the multi-view, multi-phase approach adopted by AppPoet outputs all the text successfully, and all the text is formatted, parsed and embedding, which obtains a detection accuracy of 95.50% of the ACC and 95.37% of the F1 value. In contrast, neither the multi-view-only, nor the multi-phase-only prompt engineering approach is capable of outputting the full amount of textual information. This is mainly because our experiments are based on utilizing OpenAI’s API to communicate with a remote server network. In the case of batch calls, if the prompt of a single input is too long or too much output is requested, there is a certain probability that the interaction fails due to the network problem of not being able to deliver the message properly. More seriously, the no-phase approach, due to the excessive requirements on the content of a single output from the LLM, can easily lead to the inability of the LLM to output the content in strict accordance with the preset format and requirements, as well as the inability of extracting the effective information to be transformed into embedding vectors for the discriminative process. The results of this method have a large amount of output text that cannot be utilized and transformed. In addition, the detection performance ultimately obtained by these two methods lags significantly behind that used by AppPoet. In summary, the prompt engineering approach we designed can obtain better detection performance while ensuring correct output and transformation.
Selection experiments about the number $k$ of examples. In function description generation process, we set $k$ examples to guide the LLM in producing more accurate and concise function descriptions (as shown in Table 2). To determine the optimal value of $k$ that balances effectiveness and efficiency, we conduct the experiments shown in Table 11 based on the 4000 samples randomly selected in the previous dataset.
From the results, we can draw the following conclusions. (1) In terms of both detection performance and efficiency, providing examples yields better results than not providing them, demonstrating the necessity and superiority of in-context learning. (2) The results also show that increasing the value of $k$ does not significantly improve detection accuracy but does lead to higher token and time consumption. We believe this is because the task of generating function descriptions for the relevant features is not particularly challenging for the LLM, a few examples are sufficient to clarify the desired output style and format. Based on these findings, we ultimately selected $k=3$ as the optimal number of examples for the function description prompt template.
Table 11: Selection of the number $k$ of examples.
The number of examples Feature type Function description View summary ACC F1
Token consumption Time consumption Token consumption Time consumption
Prompt Response Prompt Response
$k=0$ Permission 56.75 85.74 4.10s 2086.10 324.80 13.19s 93.75% 93.65%
API 2.78 7.32 0.71s 2214.45 381.81 15.38s
URL 4.45 13.94 1.25s 805.91 156.39 7.03s
uses-feature 0.68 1.78 0.15s
Total 64.66 108.78 6.21s 5106.46 863.00 35.6s
$k=3$ Permission 72.21 7.65 0.72s 840.08 268.18 8.64s 95.50% 95.37%
API 8.07 0.33 0.07s 728.81 249.39 7.93s
URL 12.72 0.66 0.11s 600.47 149.85 5.79s
uses-feature 2.13 0.18 0.02s
Total 95.13 8.82 0.92s 2169.36 667.42 22.36s
$k=6$ Permission 81.61 9.34 0.91s 828.99 269.29 8.61s 95.13% 95.02%
API 10.53 0.42 0.08s 732.46 252.21 8.04s
URL 15.49 0.59 0.11s 599.36 153.57 5.76s
uses-feature 2.73 0.17 0.04s
Total 110.36 10.52 1.14s 2160.81 675.07 22.41s
$k=9$ Permission 98.55 8.38 1.04s 835.82 268.04 8.92s 95.88% 95.71%
API 15.70 0.35 0.08s 736.21 253.13 8.00s
URL 17.31 0.50 0.12s 599.47 151.33 5.93s
uses-feature 3.72 0.18 0.04s
Total 135.28 9.41 1.28s 2171.5 672.5 22.85s
@@
Memory and efficiency experiments. In order to evaluate the performance of our designed memory component, we perform a set of comparison experiments. Based on the 4,000 samples randomly selected in the previous dataset, we conduct two sets of experiments with memory and without memory respectively while the multi-view text generator is working, and count the average prompt token consumption and response token consumption of each application in the experiments, as well as the the average time taken to generate text. The difference between these two sets of experiments is only whether the memory component is used to store functions. The results are shown in Fig 3.